5710 Fall 2013. Home Work assignment 1

August 28, 2013

2. Consider a quadratic form x^{T} AX where

A = ( \begin{matrix}1 & 1 & 0 & 0 \\ 1 & 1 & 0 & 0 \\ 1 & 1 & 0 & 0 \\ 1 & 1 & 0 & 0 \end{matrix} )

1